OpenStudy (anonymous):

Can anyone show me how to do this problem and explain how do do it? [2x 3 3z]=[5 3y 9]

OpenStudy (anonymous):

im sorry buti dont understand the problem even though im really good at math sorry i cant help you unless you give me more detail

OpenStudy (anonymous):

Solve The Equation haha Sorry forgot to put tht. :)

OpenStudy (anonymous):

do you know what z and y are and is x a symbol also or is it a multiplication sign

OpenStudy (anonymous):

like what they stand for

OpenStudy (preetha):

You have three unknowns, you need three equations at least to solve it.

OpenStudy (anonymous):

thts what i was thinking preetha

OpenStudy (anonymous):

the book does not say it says: Solve each equation: 1). [2x 3 3z]=[5 3y 9] thats what it says to do

OpenStudy (anonymous):

= stands for multiply rite?

OpenStudy (anonymous):

no equal means that they are gonna equal the same amount when done

OpenStudy (anonymous):

oh. idk how we could do tht then

OpenStudy (anonymous):

so what u basically have to do is think of a number for x and z and then solve the equation and think of a number for y so that when the equation y is solved in it will equal the same answer as u got when u solved the equation with x and z in it

OpenStudy (anonymous):

i found an example in the book.. want me to post it?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\left(\begin{matrix}y \\ 3x\end{matrix}\right)=\left[\begin{matrix}6 - 2x \\ 31 + 4y\end{matrix}\right]\] y=6-2x 3x=31+4y This system could be solved by using substitution: 3x=31+4y Second Equation 3x=31+4(6-2x) Substitute 6-2y for y 11x=55 Add 8x to each side x=5 Divide each side by 11 To find the value for y, substitute 5 for x in either equation: y=6-2x First equation y=6-2(5) Substitute 5 for x y=-4 Simplify The Solution is (5,-4)

OpenStudy (anonymous):

() is for matrices

OpenStudy (anonymous):

yea i dont understand this sorry

OpenStudy (anonymous):

ok

OpenStudy (preetha):

Ok, that makes sense. This is matrix algebra. Format your question in a matrix form. And substitute the first line in the second, just like the example.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

soo i would do it that to this one?? [2x 3 3z]=[5 3y 9] but if soo i dont see how, can you show me preetha?